OpenStudy (anonymous):

Quick Computing Company produces calculators. They have found that the cost, c(x), of making x calculators is a quadratic function in terms of x. The company also discovered that it costs $61 to produce 3 calculators, $165 to produce 7 calculators, and $385 to produce 12 calculators. Find the total cost of producing 6 calculators

OpenStudy (callisto):

A similar question: http://openstudy.com/study#/updates/55573d0ee4b0b98f70b37242 As for your question, since the function is quadratic in x, we have c(x) = ax^2 + bx + c Also, we are given when 3 calculators are produced, the cost is $61, i.e. c(3) = 61. When 7 calculators are produced, the cost is $165,i.e. c(7) = 165 and when 12 calculators are produced, the cost is $385, i.e. c(12)=385. So, we have the system: 61 = (3^2) a + 3b + c 165 = (7^2) a + 7b + c 385 = (12^2 ) a + 12b + c Now, you need to simplify the equations and solve the system. After you solve the system, you have the values of a, b and c. Then you can find c(6).

OpenStudy (anonymous):

I'm still confused on the last part. I'm sorry I'm difficult!

OpenStudy (callisto):

Which part?